The formula to find the area of a triangle is: Area = 1/2 the base times the height or A = 1/2BH. If you know the area of a triangle = 60 sq ft and the base = 6 ft, what's the height?

Answer by checkley71(8403)

60=6h/2
120=6h
h=120/6
h=20 answer for the height.
proof:
60=6*20/2
60=120/2
60=60